A person driving her car at 20m/s spots at a distance of 40m ahead a deer trying to cross the road. (a) If the maximum acceleration of the car is -6m/s2 , which is the maximum reaction time Δt of the driver so that the car avoids hitting the deer? (b) If the drivers reaction time is 1.2s what is the speed with which the car will hit the deer? (c) Sketch a velocity-time graph for the car.


1
Expert's answer
2021-10-22T08:29:13-0400
"40=20t+\frac{20^2}{2(40-20t)}\\t=1.3\ s"

"v=0\frac{m}{s}"